# Integrable open elliptic Toda chain with boundaries

### v1

@@ 4 XYZ chain with boundaries
Following [13] consider the monodromy matrix:
@@
due to (16), and $K^{\pm}(z)$ are $2\times 2$ matrices solving the classical quadratic reflection equation
@@
with the elliptic $r$-matrix from (16). Then the classical transfer-matrix ${\rm tr}{\mathcal{T}}^{\hbox{\tiny{openXYZ}}}(z)$ is a generating function of commuting Hamiltonians since
@@
Notice that $({{\mathcal{L}}}^{a})^{-1}(-z)={{\mathcal{L}}}^{a}(z)/\det({{\mathcal{L}}}^{a}(z))$ since the functions $\varphi_{k}(z)$ (12) are odd in $z$. On the other hand, $\det{{\mathcal{L}}}^{a}(z)={\bf C}_{2}^{a}-\wp(z){\bf C}_{1}^{a}=1$ due to (17). Therefore, ${\mathcal{T}}^{-1}(-z)=(-1)^{n}{\mathcal{L}}^{n}(z){\mathcal{L}}^{n-1}(z)\dots\mathcal{L}^{1}(z)\,.$ Finally,
@@ 5 Gauge transformed $K$-matrices
$\begin{array}[]{c}\displaystyle{{\mathcal{T}}^{\hbox{\tiny{openXYZ}}}(z)=K^{+}(z)g(z,{\bf{q}}_{1})e^{P^{1}/2c}g^{-1}(z,{\bf{q}}_{1})\dots g(z,{\bf{q}}_{n})e^{P^{n}/2c}g^{-1}(z,{\bf{q}}_{n})\times}\\ \\ \displaystyle{\times K^{-}(z)g(z,{\bf{q}}_{n})e^{P^{n}/2c}g^{-1}(z,{\bf{q}}_{n})\dots g(z,{\bf{q}}_{1})e^{P^{1}/2c}g^{-1}(z,{\bf{q}}_{1})\,.}\end{array}$ (23)
Consider the gauge transformed monodromy matrix ${\mathcal{T}}^{\hbox{\tiny{openToda}}}(z)=g^{-1}(z,{\bf{q}}_{1}){\mathcal{T}}^{\hbox{\tiny{openXYZ}}}(z)g(z,{\bf{q}}_{1})$:
$\begin{array}[]{c}\displaystyle{{\mathcal{T}}^{\hbox{\tiny{openToda}}}(z)={\tilde{K}}^{+}(z)e^{P^{1}/2c}L^{[1,2]}(z)\dots L^{[n-1,n]}(z){\tilde{K}}^{-}(z)e^{P^{n}/2c}L^{[n,n-1]}(z)\dots L^{[2,1]}(z)\,,}\end{array}$ (24)
where the notation (3) is used and
$\begin{array}[]{c}\displaystyle{{\tilde{K}}^{+}(z)=g^{-1}(z,{\bf{q}}_{1}){K}^{+}(z)g(z,{\bf{q}}_{1})\,,\qquad{\tilde{K}}^{-}(z)=g^{-1}(z,{\bf{q}}_{n}){K}^{-}(z)g(z,{\bf{q}}_{n})\,.}\end{array}$ (25)
The elliptic $K$-matrices solving the reflection equation (20) are known [5]. We write them in the following form:
$\begin{array}[]{c}\displaystyle{K^{\pm}(z)=\sigma_{0}\nu_{0}+\sum\limits_{\alpha=1}^{3}\frac{\nu^{\pm}_{\alpha}}{c_{\alpha}(\tau)}\,\frac{1}{\varphi_{\alpha}(z)}\,\sigma_{\alpha}\,,\qquad c_{k}(\tau)=\Big(\frac{\theta_{5-k}(0)}{\theta_{1}^{\prime}(0)}\Big)^{2}\,,\quad k=1,2,3\,.}\end{array}$ (26)
Equivalently,
@@
In order to compute the gauged transformed $K$-matrices (25) one should use identities with doubled nodular parameter $2\tau$:
@@
In particular, it provides $\det g(z,{\bf{q}}_{1})=-\theta_{1}(z)\theta_{1}(2{\bf{q}}_{1})$. Direct computation yields
$\begin{array}[]{c}\displaystyle{{\tilde{K}}^{+}(z)=\left(\begin{array}[]{cc}{\nu^{+}_{0}}&{0}\\ {0}&{\nu^{+}_{0}}\end{array}\right)+\sum\limits_{k=1}^{3}\nu^{+}_{k}\left(\begin{array}[]{cc}{-\varphi_{k}(2{\bf{q}}_{1})}&{\varphi_{k}(2{\bf{q}}_{1},z+\omega_{k})}\\ {\varphi_{k}(-2{\bf{q}}_{1},z+\omega_{k})}&{\varphi_{k}(2{\bf{q}}_{1})}\end{array}\right)\,,}\end{array}$ (29)
where $\varphi_{k}(z,x+\omega_{k})=\frac{\theta_{1}^{\prime}(0)\theta_{5-k}(z+x)}{\theta_{1}(z)\theta_{5-k}(x)}$ for $k=1,2,3$, that is $\varphi_{k}(z,x+\omega_{k})|_{x=0}=\varphi_{k}(z)$. The expression for ${\tilde{K}}^{-}(z)$ is the same as (29) with $\nu^{+}_{k}$ being replaced with $\nu^{-}_{k}$ and ${\bf{q}}_{1}$ being replaced with ${\bf{q}}_{n}$.
@@ 6 Open elliptic Toda chain with boundaries
Let us compute the Hamiltonian for the open elliptic Toda chain generated by ${\rm tr}{\mathcal{T}}^{\hbox{\tiny{openToda}}}(z)$ (24). Namely, consider the coefficient behind $1/z^{2n-2}$ in the expansion of ${\rm tr}{\mathcal{T}}^{\hbox{\tiny{openToda}}}(z)$ near $z=0$:
@@
Notice that matrices $L^{[a,b]}(z)$ (4) have simple poles at $z=0$, while the $K$-matrices ${\tilde{K}}^{\pm}(z)$ (29) are regular at $z=0$. Therefore,
$\begin{array}[]{c}\displaystyle{H={\rm tr}\Big({\tilde{K}}^{+}(0)e^{P^{1}/2c}A^{[1,2]}\dots A^{[n-1,n]}{\tilde{K}}^{-}(0)e^{P^{n}/2c}A^{[n,n-1]}\dots A^{[2,1]}\Big)\,,\quad A^{[k,m]}=\mathop{\hbox{Res}}\limits\limits_{z=0}L^{[k,m]}(z)\,.}\end{array}$ (32)
The residues $A^{[k,m]}$ are easily calculated using the property $\mathop{\hbox{Res}}\limits\limits_{z=0}\phi(z,u)=1$:
$\begin{array}[]{c}\displaystyle{A^{[k,m]}=\left(\begin{array}[]{c}1\\ 1\end{array}\right)\otimes(b_{1}^{[k,m]}\ b_{2}^{[k,m]})\,.}\end{array}$ (33)
It is also useful to mention that
$\begin{array}[]{c}\displaystyle{{\tilde{K}}^{+}(0)=\left(\begin{array}[]{cc}{\nu^{+}_{0}}&{0}\\ {0}&{\nu^{+}_{0}}\end{array}\right)-{\tilde{f}}^{+}({\bf{q}}_{1})\left(\begin{array}[]{c}1\\ 1\end{array}\right)\otimes(1\ -1)\,,\quad{\tilde{f}}^{+}({\bf{q}}_{1})=\sum\limits_{k=1}^{3}\nu^{+}_{k}\varphi_{k}(2{\bf{q}}_{1})}\end{array}$ (34)
@@
$\begin{array}[]{c}\displaystyle{{\tilde{K}}^{-}(0)=\left(\begin{array}[]{cc}{\nu^{+}_{0}}&{0}\\ {0}&{\nu^{-}_{0}}\end{array}\right)-{\tilde{f}}^{-}({\bf{q}}_{n})\left(\begin{array}[]{c}1\\ 1\end{array}\right)\otimes(1\ -1)\,,\quad{\tilde{f}}^{-}({\bf{q}}_{n})=\sum\limits_{k=1}^{3}\nu^{-}_{k}\varphi_{k}(2{\bf{q}}_{n})\,.}\end{array}$ (35)
In particular, this yields
$\begin{array}[]{c}\displaystyle{(b_{1}^{[2,1]}\ b_{2}^{[2,1]}){\tilde{K}}^{+}(0)e^{P^{1}/2c}\left(\begin{array}[]{c}1\\ 1\end{array}\right)=2\frac{\theta_{1}({\bf{q}}_{1}-{\bf{q}}_{2})\theta_{1}({\bf{q}}_{1}+{\bf{q}}_{2})}{\theta_{1}^{\prime}(0)\theta_{1}(2{\bf{q}}_{1})}\sinh\big(\frac{{\bf{p}}_{1}}{2c}\big)\Big(2\nu_{0}^{+}\cosh\big(\frac{{\bf{p}}_{1}}{2c}\big)-{\tilde{f}}^{+}({\bf{q}}_{1})\Big)}\end{array}$ (36)
One may use an alternatively expression for ${\tilde{f}}^{+}({\bf{q}}_{1})$ (and similarly for ${\tilde{f}}^{-}({\bf{q}}_{n})$) coming from identity
$\begin{array}[]{c}\displaystyle{\big({\tilde{f}}^{+}({\bf{q}}_{1})\big)^{2}=\sum\limits_{k=0}^{3}({\bar{\nu}}_{k}^{+})^{2}\wp({\bf{q}}_{1}+\omega_{k})-\sum\limits_{k=1}^{3}({\nu}_{k}^{+})^{2}\wp(\omega_{k})\,,}\end{array}$ (37)
@@
[…]
Finally, we come to the following Hamiltonian:

### v3

@@ 4 XYZ chain with boundaries
Let us recall main steps for construction of integrable chains with boundaries following [16]. Consider the monodromy matrix:
@@
due to each Lax matrix ${\mathcal{L}}^{a}(z)$ satisfies the same quadratic relation (29), and $K^{\pm}(z)$ are $2\times 2$ matrices solving the classical quadratic reflection equation
@@
with the same elliptic $r$-matrix from (29). Then the classical transfer-matrix ${\rm tr}{\mathcal{T}}^{\hbox{\tiny{openXYZ}}}(z)$ is a generating function of Poisson commuting Hamiltonians since
@@
In order to compute ${\mathcal{T}}^{-1}(-z)=(-1)^{n}({\mathcal{L}}^{n}(-z))^{-1}({\mathcal{L}}^{n-1}(-z))^{-1}\dots({\mathcal{L}}^{1}(-z))^{-1}$ (17) entering (36) we need to find $({{\mathcal{L}}}^{a}(-z))^{-1}$. For any matrix of the form (19) we have
$\begin{array}[]{c}\displaystyle{\Big({{\mathcal{L}}}^{a}(z)\Big)^{-1}=\frac{1}{\det({{\mathcal{L}}}^{a}(z))}\Big(\sigma_{0}{\bf S}^{a}_{0}-\sum\limits_{k=1}^{3}\sigma_{k}\varphi_{k}(z){\bf S}^{a}_{k}\Big)\,.}\end{array}$ (40)
The functions $\varphi_{k}(z)$ (21) are odd in $z$:
$\begin{array}[]{c}\displaystyle{\varphi_{k}(-z)=-\varphi_{k}(z)\,.}\end{array}$ (41)
Therefore,
$\begin{array}[]{c}\displaystyle{\Big({{\mathcal{L}}}^{a}(-z)\Big)^{-1}=\frac{1}{\det({{\mathcal{L}}}^{a}(-z))}\Big(\sigma_{0}{\bf S}^{a}_{0}+\sum\limits_{k=1}^{3}\sigma_{k}\varphi_{k}(z){\bf S}^{a}_{k}\Big)={{\mathcal{L}}}^{a}(z)\,,}\end{array}$ (42)
where we also used $\det({{\mathcal{L}}}^{a}(-z))=\det({{\mathcal{L}}}^{a}(z))=1$ because of (35) in our case. Thus,
$\begin{array}[]{c}\displaystyle{{\mathcal{T}}^{-1}(-z)=(-1)^{n}{\mathcal{L}}^{n}(z){\mathcal{L}}^{n-1}(z)\dots\mathcal{L}^{1}(z)\,,}\end{array}$ (43)
and
@@ 5 Gauge transformed $K$-matrices
$\begin{array}[]{c}\displaystyle{{\mathcal{T}}^{\hbox{\tiny{openXYZ}}}(z)=}\\ \\ \displaystyle{=K^{+}(z)g(z,{\bf{q}}_{1})e^{P^{1}/2c}g^{-1}(z,{\bf{q}}_{1})\dots g(z,{\bf{q}}_{n})e^{P^{n}/2c}g^{-1}(z,{\bf{q}}_{n})\times}\\ \\ \displaystyle{\times K^{-}(z)g(z,{\bf{q}}_{n})e^{P^{n}/2c}g^{-1}(z,{\bf{q}}_{n})\dots g(z,{\bf{q}}_{1})e^{P^{1}/2c}g^{-1}(z,{\bf{q}}_{1})\,.}\end{array}$ (45)
Consider the gauge transformed monodromy matrix:
$\begin{array}[]{c}\displaystyle{{\mathcal{T}}^{\hbox{\tiny{openToda}}}(z)=g^{-1}(z,{\bf{q}}_{1}){\mathcal{T}}^{\hbox{\tiny{openXYZ}}}(z)g(z,{\bf{q}}_{1})\,.}\end{array}$ (46)
Using the notation (3) it is written in the form:
$\begin{array}[]{c}\displaystyle{{\mathcal{T}}^{\hbox{\tiny{openToda}}}(z)=}\\ \\ \displaystyle{={\tilde{K}}^{+}(z)e^{P^{1}/2c}L^{[1,2]}(z)\dots L^{[n-1,n]}(z){\tilde{K}}^{-}(z)e^{P^{n}/2c}L^{[n,n-1]}(z)\dots L^{[2,1]}(z)\,,}\end{array}$ (47)
where
$\begin{array}[]{c}\displaystyle{{\tilde{K}}^{+}(z)=g^{-1}(z,{\bf{q}}_{1}){K}^{+}(z)g(z,{\bf{q}}_{1})\,,}\\ \\ \displaystyle{{\tilde{K}}^{-}(z)=g^{-1}(z,{\bf{q}}_{n}){K}^{-}(z)g(z,{\bf{q}}_{n})\,.}\end{array}$ (48)
The elliptic $K$-matrix solving the reflection equation (38) is known from [7]. We write it in the following form:
$\begin{array}[]{c}\displaystyle{K^{\pm}(z)=\sigma_{0}\nu_{0}^{\pm}+\sum\limits_{\alpha=1}^{3}\frac{\nu^{\pm}_{\alpha}}{c_{\alpha}(\tau)}\,\frac{1}{\varphi_{\alpha}(z)}\,\sigma_{\alpha}}\end{array}$ (49)
with the set of theta-constants
$\begin{array}[]{c}\displaystyle{c_{k}(\tau)=\Big(\frac{\theta_{5-k}(0)}{\theta_{1}^{\prime}(0)}\Big)^{2}\,,\quad k=1,2,3\,.}\end{array}$ (50)
Each $K$-matrix $K^{+}(z)$ and $K^{-}(z)$ depends on four coupling constants $\nu_{i}^{+}$ and $\nu^{-}_{i}$, $i=0,1,2,3$ respectively. Equivalently, expressions (49) are represented as follows:
@@
Proposition The gauge transformed $K$-matrices ${\tilde{K}}(z)$ (48) defined through $K^{\pm}(z)$ (49) with the gauge transformation (2) take the following form:
$\begin{array}[]{c}\displaystyle{{\tilde{K}}^{+}(z)=\left(\begin{array}[]{cc}{\nu^{+}_{0}}&{0}\\ \\ {0}&{\nu^{+}_{0}}\end{array}\right)+\sum\limits_{k=1}^{3}\nu^{+}_{k}\left(\begin{array}[]{cc}{-\varphi_{k}(2{\bf{q}}_{1})}&{\varphi_{k}(2{\bf{q}}_{1},z+\omega_{k})}\\ \\ {\varphi_{k}(-2{\bf{q}}_{1},z+\omega_{k})}&{\varphi_{k}(2{\bf{q}}_{1})}\end{array}\right)\,,}\\ \\ \displaystyle{{\tilde{K}}^{-}(z)=\left(\begin{array}[]{cc}{\nu^{-}_{0}}&{0}\\ \\ {0}&{\nu^{-}_{0}}\end{array}\right)+\sum\limits_{k=1}^{3}\nu^{-}_{k}\left(\begin{array}[]{cc}{-\varphi_{k}(2{\bf{q}}_{n})}&{\varphi_{k}(2{\bf{q}}_{n},z+\omega_{k})}\\ \\ {\varphi_{k}(-2{\bf{q}}_{n},z+\omega_{k})}&{\varphi_{k}(2{\bf{q}}_{n})}\end{array}\right)\,,}\end{array}$ (52)
where the set of functions
$\begin{array}[]{c}\displaystyle{\varphi_{k}(z,x+\omega_{k})=\frac{\theta_{1}^{\prime}(0)\theta_{5-k}(z+x)}{\theta_{1}(z)\theta_{5-k}(x)}\,,\quad k=1,2,3\,,}\end{array}$ (53)
$\begin{array}[]{c}\displaystyle{\varphi_{k}(z,x+\omega_{k})\big|_{x=0}=\varphi_{k}(z)\,.}\end{array}$ (54)
is used.
The proof is by direct computation using identities with doubled modular parameter $2\tau$:
@@
In particular, it provides $\det g(z,{\bf{q}}_{a})=-\theta_{1}(z)\theta_{1}(2{\bf{q}}_{a})$.
Let us remark that similar calculation was performed for relation between ${\rm BC}_{1}$ Calogero-Inozemtsev model and the non-relativistic Zhukovsky-Volterra gyrostat in [10] as well as for relation between ${\rm BC}_{1}$ Ruijsenaars- van Diejen model and the relativistic Zhukovsky-Volterra gyrostat in recent paper [11].
@@ 6 Open elliptic Toda chain with boundaries
Let us compute the Hamiltonian for the open elliptic Toda chain generated by ${\rm tr}{\mathcal{T}}^{\hbox{\tiny{openToda}}}(z)$ (47).
Transfer-matrix.
Consider the coefficient behind $1/z^{2n-2}$ in the expansion of ${\rm tr}{\mathcal{T}}^{\hbox{\tiny{openToda}}}(z)$ near $z=0$:
@@
Notice that matrices $L^{[a,b]}(z)$ (5) have only simple poles at $z=0$ due to
$\begin{array}[]{c}\displaystyle{\mathop{\hbox{Res}}\limits\limits_{z=0}\phi(z,u)=1}\end{array}$ (58)
for the function $\phi(z,u)$ (6). The product ${\rm tr}{\mathcal{T}}^{\hbox{\tiny{openToda}}}(z)$ (47) contains 2n22n-2 matrices $L^{[a,b]}(z)$, and the boundary $K$-matrices ${\tilde{K}}^{\pm}(z)$ (52) are regular at $z=0$. Therefore,
$\begin{array}[]{c}\displaystyle{H={\rm tr}\Big({\tilde{K}}^{+}(0)e^{P^{1}/2c}A^{[1,2]}\dots A^{[n-1,n]}{\tilde{K}}^{-}(0)e^{P^{n}/2c}A^{[n,n-1]}\dots A^{[2,1]}\Big)}\end{array}$ (59)
with
$\begin{array}[]{c}\displaystyle{A^{[k,m]}=\mathop{\hbox{Res}}\limits\limits_{z=0}L^{[k,m]}(z)\,.}\end{array}$ (60)
The residues $A^{[k,m]}$ are easily calculated using the property (58):
$\begin{array}[]{c}\displaystyle{A^{[k,m]}=\left(\begin{array}[]{cc}{b_{1}^{[k,m]}}&{b_{2}^{[k,m]}}\\ {b_{1}^{[k,m]}}&{b_{2}^{[k,m]}}\end{array}\right)=\left(\begin{array}[]{c}1\\ 1\end{array}\right)\otimes(b_{1}^{[k,m]}\,,b_{2}^{[k,m]})\,,}\end{array}$ (61)
where $(b_{1}^{[k,m]}\,,b_{2}^{[k,m]})$ are 2-dimensional row-vectors.
Expressions for ${\tilde{K}}^{\pm}(0)$.
Taking into account the oddness (41) of $\varphi_{k}(z)$ and the definitions (53)-(54), from (52) we have
$\begin{array}[]{c}\displaystyle{{\tilde{K}}^{+}(0)=\left(\begin{array}[]{cc}{\nu^{+}_{0}}&{0}\\ \\ {0}&{\nu^{+}_{0}}\end{array}\right)+\sum\limits_{k=1}^{3}\nu^{+}_{k}\left(\begin{array}[]{cc}{-\varphi_{k}(2{\bf{q}}_{1})}&{\varphi_{k}(2{\bf{q}}_{1})}\\ \\ {-\varphi_{k}(2{\bf{q}}_{1})}&{\varphi_{k}(2{\bf{q}}_{1})}\end{array}\right)\,,}\\ \\ \displaystyle{{\tilde{K}}^{-}(0)=\left(\begin{array}[]{cc}{\nu^{-}_{0}}&{0}\\ \\ {0}&{\nu^{-}_{0}}\end{array}\right)+\sum\limits_{k=1}^{3}\nu^{-}_{k}\left(\begin{array}[]{cc}{-\varphi_{k}(2{\bf{q}}_{n})}&{\varphi_{k}(2{\bf{q}}_{n})}\\ \\ {-\varphi_{k}(2{\bf{q}}_{n})}&{\varphi_{k}(2{\bf{q}}_{n})}\end{array}\right)\,,}\end{array}$ (62)
Put it differently,
$\begin{array}[]{c}\displaystyle{{\tilde{K}}^{+}(0)=\left(\begin{array}[]{cc}{\nu^{+}_{0}}&{0}\\ {0}&{\nu^{+}_{0}}\end{array}\right)-{\tilde{f}}^{+}({\bf{q}}_{1})\left(\begin{array}[]{c}1\\ 1\end{array}\right)\otimes(1\,,-1)}\end{array}$ (63)
with
$\begin{array}[]{c}\displaystyle{{\tilde{f}}^{+}({\bf{q}}_{1})=\sum\limits_{k=1}^{3}\nu^{+}_{k}\varphi_{k}(2{\bf{q}}_{1})}\end{array}$ (64)
@@
$\begin{array}[]{c}\displaystyle{{\tilde{K}}^{-}(0)=\left(\begin{array}[]{cc}{\nu^{-}_{0}}&{0}\\ {0}&{\nu^{-}_{0}}\end{array}\right)-{\tilde{f}}^{-}({\bf{q}}_{n})\left(\begin{array}[]{c}1\\ 1\end{array}\right)\otimes(1\,,-1)\,,}\end{array}$ (65)
$\begin{array}[]{c}\displaystyle{{\tilde{f}}^{-}({\bf{q}}_{n})=\sum\limits_{k=1}^{3}\nu^{-}_{k}\varphi_{k}(2{\bf{q}}_{n})\,.}\end{array}$ (66)
One may use an alternative representation for the functions ${\tilde{f}}^{+}({\bf{q}}_{1})$ and ${\tilde{f}}^{-}({\bf{q}}_{n})$:
$\begin{array}[]{c}\displaystyle{\big({\tilde{f}}^{+}({\bf{q}}_{1})\big)^{2}=\sum\limits_{k=0}^{3}({\bar{\nu}}_{k}^{+})^{2}\wp({\bf{q}}_{1}+\omega_{k})-\sum\limits_{k=1}^{3}({\nu}_{k}^{+})^{2}\wp(\omega_{k})\,,}\\ \\ \displaystyle{\big({\tilde{f}}^{-}({\bf{q}}_{n})\big)^{2}=\sum\limits_{k=0}^{3}({\bar{\nu}}_{k}^{-})^{2}\wp({\bf{q}}_{n}+\omega_{k})-\sum\limits_{k=1}^{3}({\nu}_{k}^{-})^{2}\wp(\omega_{k})\,,}\end{array}$ (67)
@@
$\begin{array}[]{c}\left(\begin{array}[]{c}\bar{\nu}^{\pm}_{0}\\ \bar{\nu}^{\pm}_{1}\\ \bar{\nu}^{\pm}_{2}\\ \bar{\nu}^{\pm}_{3}\end{array}\right)=\displaystyle{\frac{1}{2}}\left(\begin{array}[]{ccc}1&1&1\\ 1&-1&-1\\ -1&1&-1\\ -1&-1&1\end{array}\right)\left(\begin{array}[]{c}\nu^{\pm}_{1}\\ \nu^{\pm}_{2}\\ \nu^{\pm}_{3}\end{array}\right)\,.\end{array}$ (68)
This type identities can be found in [11].
The Hamiltonian.
Let us compute the expression for $H$ (59). Plugging (61) into (59) we get:
$\begin{array}[]{c}\displaystyle{H=\prod\limits_{a=2}^{n-1}\Big(b_{1}^{[a-1,a]}+b_{2}^{[a-1,a]}\Big)\Big(b_{1}^{[a+1,a]}+b_{2}^{[a+1,a]}\Big)\times}\\ \\ \displaystyle{\times\left[\big(b_{1}^{[2,1]}\,,b_{2}^{[2,1]}\big){\tilde{K}}^{+}(0)e^{P^{1}/2c}\left(\begin{array}[]{c}1\\ 1\end{array}\right)\right]\times\left[\big(b_{1}^{[n-1,n]}\,,b_{2}^{[n-1,n]}\big){\tilde{K}}^{-}(0)e^{P^{n}/2c}\left(\begin{array}[]{c}1\\ 1\end{array}\right)\right]\,.}\end{array}$ (69)
It follows from (7) that
$\begin{array}[]{c}\displaystyle{b_{1}^{[a-1,a]}+b_{2}^{[a-1,a]}=2\frac{\theta_{1}({\bf{q}}_{a}-{\bf{q}}_{a-1})\theta_{1}({\bf{q}}_{a}+{\bf{q}}_{a-1})}{\theta_{1}^{\prime}(0)\theta_{1}(2{\bf{q}}_{a})}\sinh\big(\frac{{\bf{p}}_{a}}{2c}\big)\,,}\\ \\ \displaystyle{b_{1}^{[a+1,a]}+b_{2}^{[a+1,a]}=2\frac{\theta_{1}({\bf{q}}_{a}-{\bf{q}}_{a+1})\theta_{1}({\bf{q}}_{a}+{\bf{q}}_{a+1})}{\theta_{1}^{\prime}(0)\theta_{1}(2{\bf{q}}_{a})}\sinh\big(\frac{{\bf{p}}_{a}}{2c}\big)\,.}\end{array}$ (70)
Then for the upper line of (69) one obtains:
$\begin{array}[]{c}\displaystyle{\prod\limits_{a=2}^{n-1}\Big(b_{1}^{[a-1,a]}+b_{2}^{[a-1,a]}\Big)\Big(b_{1}^{[a+1,a]}+b_{2}^{[a+1,a]}\Big)=}\\ \\ \displaystyle{=\theta_{1}({\bf{q}}_{2}-{\bf{q}}_{1})\theta_{1}({\bf{q}}_{2}+{\bf{q}}_{1})\theta_{1}({\bf{q}}_{n-1}-{\bf{q}}_{n})\theta_{1}({\bf{q}}_{n-1}+{\bf{q}}_{n})\times}\\ \\ \displaystyle{\times(-1)^{n-3}\Big(\frac{2}{\theta_{1}^{\prime}(0)}\Big)^{2n-4}\,\left(\frac{\displaystyle{\prod\limits_{a=3}^{n-1}\theta_{1}({\bf{q}}_{a}-{\bf{q}}_{a-1})\theta_{1}({\bf{q}}_{a}+{\bf{q}}_{a-1})}}{\displaystyle{\prod\limits_{a=2}^{n-1}\theta_{1}(2{\bf{q}}_{a})}}\right)^{2}\,\prod\limits_{a=2}^{n-1}\sinh^{2}\big(\frac{{\bf{p}}_{a}}{2c}\big)\,.}\end{array}$ (71)
Consider the lower line of (69). Using (7) and (63) we have